What is the terminal speed for an 85 kg skier going down a 35 ∘ snow-covered slope on wooden skis μk= 0.060?Assume that the skier is 1.8 m tall and 0.50 m wide. Assume the skier's drag coefficient is 0.80.
Answer #1

At equilibrium (constant velocity), the downslope component of weight is equal to the sum of the upslope forces of friction and drag:
mgsinΘ = µmgcosΘ + ½ρCAv²
For the drag term, take ρ = 1.293 kg/m³
and C = 1
and A = 1.8m * 0.5m = 0.9 m². Then
85kg * 9.8m/s² * sin35º = 0.06 * 85kg * 9.8m/s² * cos35º
+ ½ * 1.293kg/m³ * 1 * 0.9m² * v²
478 N = 41N + 0.58kg/m * v²
v² = (478 - 41)N / 0.58kg = 753 m²/s²
v = 27.4 m/s